OpenStudy (anonymous):

While simplifying some math work, Peter wrote on his paper that x^3 • x^3 • x^3 • x^3 equaled x^3+ ^3 + ^3 +^3 . Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify x^3 + x^3 + x^3 + x^3?

OpenStudy (anonymous):

I know the formula is a^m * a^n = a ^m + ^n. Though I think that something like x^3 = x * x *x right so x^3 + ^3 + ^3 +^3 would be correct right?

OpenStudy (anonymous):

can you write exponents with no base?

OpenStudy (anonymous):

I don't think so? the base would be x right?

zepdrix (zepdrix):

issy, yes you have the right idea. x^3 • x^3 • x^3 • x^3 = x^{3+ 3 + 3 +3} ^(I'm assuming this is what it's supposed to look like?) `Did peter simplify his work correctly?` yes. `and completely?` no. It can be further simplified, do you see how?

OpenStudy (anonymous):

completely would be x^12 right?

zepdrix (zepdrix):

yah looks good :)

OpenStudy (anonymous):

and its asking if the work would be the same if it was simplified and I think it would right?

OpenStudy (anonymous):

\[x^{3+3+3+3}\] this one is equivalent so if x^3 +^3 +^3 +^3 is like this then it's a correct simplification

zepdrix (zepdrix):

The next part is something different,\[\Large x^3+x^3+x^3+x^3 \quad=\quad?\]

OpenStudy (anonymous):

x^3 +^3+^3+^3 = x^12 right?

OpenStudy (anonymous):

so his answer would be the same if he simplified it more right?

zepdrix (zepdrix):